Calculus I, Fall 2021 (Thomas’ Calculus Early Transcendentals 13ed), http://www.math.nthu.edu.tw/˜ wangwc/

Brief solutions to selected problems in homework week 04

Figure 1: Correct answer section 2.5, problem 68 (only minor mistake)

Figure 2: A common mistake of section 2.5, problem 68. One should at least state how to choose ε when f(c) < 0